The following information is for the Jeffries​ Corporation:

Product​ A: Revenue $ 15.00

Variable Cost $ 10.00

Product​ B: Revenue $ 33.00

Variable Cost $ 18.00

Total fixed costs $ 399

What is the breakeven​ point, assuming the sales mix consists of three units of Product A and one unit of Product B
Answer #1

Weighted average contribution margin per unit = (15-10*3/4+33-18/4) = 7.5

Break even unit = 399/7.5 = 53.20 Units

Product A = 53.20*3/4 = 39.90 Units

Product B = 53.20/4 = 13.30 Units